Grade 12th passPhysical Chemistry

a system absorb 640 h heat and does work of 260 j the change in internal energy of the system will be?​

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Pawan Prajapati
5 Years ago
Answer:

The system absorbs 640 J and the work system does on surroundings is 260J. Therefore by the first law of thermodynamics, the internal energy of the system increases by: △U=q−W=(640−260)=+380 J.

Thanks